O output and the factors of production. QUESTION 19 Consider the following production function: Y-k4s whens capital stock is 144, ourput s Ο 12. 24. ○ 144. O 6. QUESTION 20 What ere the factors of on
Answer #1

production function: Y=K^0.5

Y=K^0.5

Y=(144)^0.5

Y=12 is answer..

the above is answer..